If you have two parabolas formed by the two equations: y=x^2+2 and y=-x^2+5 at what point(s) do these two parabolas intersect each other?
Answer by Alan3354(69443)

If you have two parabolas formed by the two equations: y=x^2+2 and y=-x^2+5 at what point(s) do these two parabolas intersect each other?
-----------
They both equal y, so
x^2 + 2 = -x^2 + 5
2x^2 = 3
x^2 = 3/2
Solve that for x, find the y's